Terminal points P(x,y)
1
Expert's answer
2021-09-01T11:22:03-0400

Let "P(x, y)" be the terminal point on the unit circle determined by "t." Then


"\sin t=y, \cos t=x, \tan t=\dfrac{y}{x}"

"x^2+y^2=\cos^2 t+\sin^2t=1"